Texas Natural Gas has $1,000 par value bonds outstanding at 12% interest. The bonds will mature in 50 years. Compute the current price of the bonds if the yields to maturity (YTM) is 14 percent. Assume annual coupon payments
Answer #1

Annual coupon=1000*12%=120

Hence current price=Annual coupon*Present value of annuity factor(14%,50)+1000*Present value of discounting factor(14%,50)

=120*7.13265646+1000*0.0014280961

=$857.35(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=120[1-(1.14)^-50]/0.14

=120*7.13265646

2.Present value of discounting factor=1000/1.14^50

=1000*0.0014280961